OpenStudy (anonymous):

how do i solve for y in this equation: 4x-5xy=7

OpenStudy (anonymous):

dude like u cud probably do some algebra or something mang

OpenStudy (anonymous):

maybe like -4x on both sides and see wut u get

OpenStudy (anonymous):

then you could like totally divide afterwards

OpenStudy (imstuck):

Be nice...it's his first time

OpenStudy (anonymous):

?

OpenStudy (loser66):

?? from me, too. hihihi

OpenStudy (anonymous):

do you have answer choices?

OpenStudy (anonymous):

4x - 5xy = 7 -5xy = 7 - 4x 5xy = 4x - 7 y = (4x - 7)/(5x) x^2 + y = -4 y = -4 - x^2

OpenStudy (anonymous):

Simplifying 4x + -5xy = 7 Solving 4x + -5xy = 7 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Reorder the terms: -7 + 4x + -5xy = 7 + -7 Combine like terms: 7 + -7 = 0 -7 + 4x + -5xy = 0 The solution to this equation could not be determined.

OpenStudy (anonymous):

its the first on 4x-5xy+7

OpenStudy (anonymous):

I found an answer to the equation. First you'd have to subtract the 4x to both sides. Then you would divide -5xy by -5x to have y by itself. The 7 and -4x would be divided by the -5x. Then the answer would be y=-7/5x+4/5.

OpenStudy (anonymous):

@tatiana__celeste__ thats answer not apart of the answer chose im going to need u to read what was said plz and thankz